OpenStudy (anonymous):

find the derivative of y=x^2 with respect to lnx

OpenStudy (raden):

u meant : y = x^2 * ln(x) ?

OpenStudy (anonymous):

ummm im not sure thats what my hw problem says

OpenStudy (anonymous):

take z= lnx and we need to find dy/dz =(dy/dx) / (dz/dx)

OpenStudy (anonymous):

here dy/dx=2x and dz/dx =1/x hence dy/dz =2x/(1/x) =2x^2 or in other words the derivative of y=x^2 with respect to lnx is 2x^2

OpenStudy (anonymous):

ohhhh thank you

OpenStudy (anonymous):

welcome dear